OpenStudy (anonymous):

In a bag, there are 2 blue tokens, 3 red tokens, 6 yellow tokens, and 4 green tokens. Stephen will perform random trials of drawing a token and replacing it into the bag. Which of the following is a good prediction for the number of times a red or green token will be drawn in 300 trials? it is hard

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

thanks for helping me :D

OpenStudy (anonymous):

P(red or green) = (3+4)/(15) = 7/15 so in 300, you would expect 7/15 (300) = 140 (red or green)

OpenStudy (anonymous):

thanks XD

OpenStudy (anonymous):

:)